Status Report

ISS On-Orbit Status 9 Mar 2002

By SpaceRef Editor
March 9, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted
previously or below. First of two weekend rest days.

After wake-up at 1:00 am EST, Carl Walz and Dan Bursch received
compliments and thanks from Flight Control for their perseverance
yesterday in replacing the BCM (battery charger module) in the
Airlock (A/L) despite tight clearances of the A/L rack.

As “over” at STS-109 the upgraded and vastly more powerful Hubble
Space Telescope was successfully set free in space from the Columbia
PLB at 5:04 am EST, the ISS crew was busy performing the regular
weekly 2-hr. house cleaning.

Crew was also thanked for their help with the VRS (vacuum resource
system) troubleshooting tests. Results of the tests show excessive
leakage at the mated connection of the vacuum QD (quick disconnect)
at the Rack/line interface. MCC-H will decide later whether or
not to use the QD port “as is”. In the meantime, the HRF (human
research facility) does not require use of this particular vacuum

SD2 (smoke detector #2) in the Node has been inhibited for some time
and is not providing usable data. The ground has decided to
command its RPC (remote power controller) Open, since there is a
concern about the integrity of the cabling in the wire harness.
SD1 continues to operate normally and provides adequate smoke
detection for the Node.

At 2:00 pm EST, FE-1 Walz began to deactivate the EarthKAM payload
and to dismantle its equipment at the Lab window, after the remotely
controlled system had been capturing Earth images all day.
EarthKAM components were stowed for later use.

Imagery of the SAMS (space acceleration measurement system) payload
was placed on the crew’s “job jar” task list. Several digital
photos are to be taken of the SAMS sensor head that is mounted near
the top of Express Rack 2 (ER2). POIC (Payload Operations
Integration Center) wants to capture views showing the orientation of
the sensing equipment.

The crew was requested to provide information on the CEVIS (cycle
ergometer with vibration isolation) to allow MCC-H to make a final
determination of whether future CEVIS refurbishment is required. The
information pertains to the long isolators, the condition of their
wires and their shapes.

The crew performed a full set of physical exercise (one hour on
treadmill or CEVIS, 1.5 hrs. on RED).

The crews of ISS and STS-109 exchanged ship-to-ship greetings today
(8:30 am EST) on S/G2 (space-to-ground #2) in a communications check
for tomorrow’s PAO TV event between ISS and Shuttle Columbia. The
ship-to-ship tomorrow morning is scheduled for 3:15 3:27am EST,
with video on Ku-band and audio on S/G2.

Target areas for the Russian Uragan (“hurricane”) earth imaging
experiment were placed on Yuri Onufrienko’s task list for today and
tomorrow, covering numerous science-related observations and
photo/video ops from SM window #9. For example: a strong earthquake
measuring 7 on the Richter scale struck Afghanistan on 3/3-4.
Imagery of the Medvezhi Glacier in the aftermath of this
earthquake is of particular interest. Using the Rubinar
binocular system, Onufrienko was to image population centers in the
desert, the Salang Pass 160 km left of the track, the city of Kabul
and surrounding area, and other sites of interest.

A new TDRS (tracking and data relay satellite) was successfully
launched last evening on an Atlas 2A: TDRS-9 (TDRS-I), supporting
S-band, Ku-band and Ka-band communications.

Science Update (Expedition Four — 13th):
As the lead scientist reported: A week filled with
accomplishments, as well as with troubleshooting and recovery
activities. ARIS appears to be close to being up and running
nominally again the ground believes to have a lead on EXPPCS, and
EarthKAM was concluded despite a bumpy start.

Hoffman-Reflex: The last in-flight session for this
Increment is scheduled for next week.

Extra-Vehicular Activity Radiation Monitors (EVARM): The
last data set downlinked looks good. Also, EVARM support was happy to
hear that the sessions today were completed ahead of schedule.

Ultrasound: In progress.

Pulmonary Function in Flight (PuFF): Files from the crew’s
PuFF sessions were successful downlinked yesterday and are being
forwarded to the PI team for analysis.

Renal (Kidney) Stone Experiment: In progress.

Interactions (NTXN): The Interactions Team was thrilled
that all questionnaires were completed on Wednesday, the target day
for data collection. Data are also collected from ground
subjects each week, likewise on Wednesdays.

Human Research Facility Workstation (HRF WS): n/a

Human Research Facility/PC (HRF/PC): n/a

Cellular Biotechnology Support Systems (CBOSS): The
CBOSS team continues to monitor the temperature in the BTR
(biotechnology refrigerator).

Physics of Colloids in Space (EXPPCS): Ground
troubleshooting continues on the EXPPCS flight system computer
boot-up problem experienced on 2/24 (Sunday) evening. The ground
teams are currently assessing a potential corrupted BIOS (basic
input/output system) configuration file. By disabling the BIOS
on the ground engineering unit and trying to boot up, the EXPPCS team
could reproduce the same power draw profile as was seen during the
2/24 on-orbit failure to boot. As a result, EXPPCS has looked in
great detail into the various BIOS settings (CMOSS, flash memory, and
factory BIOS) resident on our CPU card, and the feeling is that the
EXPPCS flight system CPU configuration settings have been corrupted
or erased. The EXPPCS team is looking at options for on-orbit

Space Acceleration Measurement System (SAMS): SAMS hardware is

Microgravity Acceleration Measurement System (MAMS): MAMS
continues collection and archival of low-frequency acceleration
data for the characterization of the ISS quasi-steady microgravity

Protein Crystal Growth-Single Locker Thermal Enclosure System
Both STES units are nominal.

Materials ISS Experiment (MISSE): Nominal and collecting

Educational Payload Operations (EPO): EPO activities are

Active Rack Isolation System – Isolation Characterization
Experiment (ARIS-ICE)
: ARIS recovery is nearly complete.
Ground-commanded checkout testing, which will complete on 3/12,
indicates everything is performing well and no problems exist.
In their hard work repairing ARIS, Carl and Dan rotated the
ARIS rack for the first time since its initial installation, a very
significant step because it gained the ground very important
knowledge that is leading to improvements in the ARIS procedures,
decreasing the risk of future ARIS hardware problems.

EarthKAM: The ground has so far received 169 images on a
total of 16 orbits. 17 of 20 participating schools have
submitted photo requests, and more continue to come on-line and begin
taking pictures. The 180-mm lens was installed on the camera
yesterday at approximately 7:20 am EST. The ground has not yet
received any images taken with this lens, but anticipates excellent
and high-resolution images similar to those from the February
operations. “The schools can’t wait!” All EarthKAM images are
available for public access on the EarthKAM data system at:

Advanced Astroculture (ADVASC): Comprehensive analysis
of ADVASC’s science/health and status data and images suggests that
the plants have been developing slower than expected. The pictures
taken by the crew yesterday (3/8) should provide a better idea of how
much the plants have developed and will help to determine when to
reschedule the plant tissue sample activities.

Zeolite Crystal Growth (ZCG): Planned. Hardware is
ready to process samples to be launched on ISS 8A.

Crew Earth Observations (CEO): Thus far, 2725 ESC images
have been downlinked by the Exp. 4 crew 920 have been screened for
content and image quality and released for posting on the STIC.
Ground has acquired 40 to 50 digital images of target research sites
and of dynamic phenomena such as volcanic eruptions researchers are
looking forward to the return of the Hasselblad 70-mm film. An
ISS 2 image of the Montserrat volcano with a plume of smoke and ash
(7/9/01) is the subject of a web tutorial on image transformation
submitted to the Earth Sciences &amp Image Analysis Lab. The author
describes the way in which this excellent oblique view was registered
to a map and transformed to a pseudo-nadir view. Among notable
digital images from Exp. 4 is one of Santa Maria volcano on the
outskirts of Quetzaltenango, Guatemala. An article featuring that
image has been published on the NASA-wide Earth Observatory

Today’s optional CEO target areas were Tigris-Euphrates,
(an excellent mapping pass with good lighting and
weather. Of interest: careful mapping of the reservoirs and lakes of
this headwater region for the Tigris-Euphrates Rivers. If time
permitted, crew was to continue photographing targets further down
the valley including Baghdad and Kuwait)
, E. Mediterranean
Dust and Smog
(a weakening storm system and cold front remain
stalled over the western and central Med. As ISS approached the coast
of north Africa, crew was to look to the right of track for dust
plumes moving northward out of Egypt)
, Somalia Coast
(with most of the target area to the right of track this pass,
crew was asked to try for oblique, context views looking
southwestward down the coast)
, W. Mediterranean Dust and
(an unusual winter storm is slowly breaking up over the
western Med. Of interest: looking to the right of track, as ISS
crossed southern Spain and eastern Morocco, for rare snow cover in
the mountains there)
, Rukwa Transform, Tanzania (only
isolated thunderstorm were expected in the vicinity of the Rukwa
Transform this pass. Of interest: mapping the details of this complex
of faults and folds with a high sun angle)
, Industrialized
Southeastern Africa
(high pressure is breaking down over South
Africa this pass. However, crew was to wait until ISS reached the
coat and then to try for oblique and limb shots to the right of track
as this air mass and its aerosols move out to sea)
, and
Eastern United States (as ISS tracked out of eastern Canada
over northern New England, crew was to look to the right of track
both sea fog and land smog patterns under the retreating high
pressure area)

ISS Orbit (as of this morning, 6:45am EST):
Mean altitude — 386.4 km

  • Apogee — 389.6 km
  • Perigee — 383.1 km
  • Period — 92.3 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0004832
  • Orbits per 24-hr. day — 15.60
  • Altitude decrease — 380 m (mean) in last 24 hours
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 18848
  • Current Flight Attitude — LVLH (local vertical/local horizontal =
  • “earth-fixed” [yaw: -10 deg, pitch: -7 deg., roll: 0 deg])roll: 0
  • deg]).

For more on ISS orbit and naked-eye visibility dates/times,

SpaceRef staff editor.